OpenStudy (anonymous):

jit4wonsolve each equation by completing the square. round to the nearest hundredth if necessary. 3. x^2+6x =16 4. x^2-10x =11 5. x^2-9x =0 6. x^2+16x =15 7. 3x^2+18x-81=0 use the quadratic formula to solve each equation. round to the nearest hundredth if necessary. 8. 4x^2+3x-8=0 9.2x^2-7x+3=0 10. x^2-2x+3=0 11. 2x^2+4x-7=0 I DONT GET THIS STUFF.

OpenStudy (anonymous):

so for quadratic stuff.. Use this formula..\[-b \pm \sqrt{b ^{2}-4ac}/2a\] like for question 8: a=4; b=3; c=-8 <-- plug these number in the equation and solve..

OpenStudy (anonymous):

i knew i was missing a formula lol. thanks.

OpenStudy (anonymous):

lol kk welcome.. for completing the sq.. http://www.purplemath.com/modules/sqrquad.htm

OpenStudy (anonymous):

for ur question 3. \[x ^{2}+6x=16\] \[x ^{2}+6x+9\] is a complete sq. so u take 3 inside the brackets and take +9 to the other side.. like, \[(x+3)^{2}=16+9\]

OpenStudy (anonymous):

\[(x+3)^{2}=25\] <-- take the sq. root on both sides.. \[\sqrt{(x+3)^{2}}=\sqrt{5^{2}}\]

OpenStudy (anonymous):

\[x+3=\pm5\]

OpenStudy (anonymous):

are these the answers?

OpenStudy (anonymous):

x=5-3 ; x=-5-3 x=2 ; x=-8 No this is answer to question 3

OpenStudy (anonymous):

oh okay thanks.

OpenStudy (anonymous):

y welcome

OpenStudy (anonymous):

can you do the rest lol?

OpenStudy (anonymous):

try doing it, if you get stuck i will help u..

OpenStudy (anonymous):

i tried and im stuck. and i have to turn these problems in in 20 mins and its a test.

OpenStudy (anonymous):

X^2-6X=20 its saying round to the nearest hundredth can someone help me please